OpenStudy (darkbluechocobo):

Identify the x-intercept(s) for: f(x) = x^2 + 3x − 4

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

lol

OpenStudy (triciaal):

x-intercepts when y= 0 f(x) = 0 factor the quadratic

OpenStudy (darkbluechocobo):

so (1,0) through (-4,0)

OpenStudy (triciaal):

already general form a = 1 b = 3 and c = -4 what is a*c what are the factors that add up to b? do you still want to go through the steps?

OpenStudy (darkbluechocobo):

a*c=-4. when you say add up do you mean multiply into b?

OpenStudy (cwrw238):

you need 2 numbers whose product is -4 and whose sum = 3

OpenStudy (darkbluechocobo):

-4 and 1

OpenStudy (cwrw238):

if 2 numbers are a and b then ab = -4 a+b = 3

OpenStudy (isaiah.feynman):

The x intercepts is also the roots of the equation.

OpenStudy (cwrw238):

swap the signs

OpenStudy (cwrw238):

-4 + 1 = -3

OpenStudy (cwrw238):

so not those 2 numbers

OpenStudy (darkbluechocobo):

so -1 and 4

OpenStudy (cwrw238):

right!

OpenStudy (darkbluechocobo):

hokay thank you for working that out with me

OpenStudy (cwrw238):

now those 2 numbers give you the factors x^2 +3x - 4 = 0 (x -4)(x +1) = 0

OpenStudy (darkbluechocobo):

question though with the graph of this that was posted earlier it was -4 and positive 1 why did this switch?

OpenStudy (cwrw238):

oh sorry - i made a mistake its (x + 4)(x - 1) = 0

OpenStudy (darkbluechocobo):

Alrights so basically the x-intercepts come from multiplying a*c and then finding what two numbers multiply into c and add into b?

OpenStudy (cwrw238):

having factored the function we now can find the values of x when f(x) is zero

OpenStudy (cwrw238):

- those 2 numbers do NOT give you the x-intercepts - they give you the values that go into the two factors

OpenStudy (darkbluechocobo):

ok

OpenStudy (darkbluechocobo):

so then what is the next step

OpenStudy (cwrw238):